At the beginning of a semester, Dirk is willing to pay $180 for a textbook, but gets it for $140 at the university bookstore. When he has completed the course, the book is worth only $40 to him. He can keep it or sell it back to the bookstore. The bookstore will pay him 35% of the original price. If Dirk makes his best choice, overall, he will (gain or lose) a value of $_.
Answer #1

Dirk surplus while purchasing the book = Willingness to pay - Actual price paid for book = 180 - 140 = $ 40

Dirk surplus while selling th book = Price paid by the bookstore - Book worth for Dirk = 35% of 140 - 40 = 49 - 40 = $ 9

So, total gain = 40 + 9 = $ 49